Shift summary for Owl, 8/3/99

Ed Blucher (blucher@hep.uchicago.edu)
Tue, 3 Aug 1999 06:52:26 -0500

Shift Summary for Owl Shift, 08/3/99

On Shift: EB,BV

Hours of beam available at all intensities: 5.5
Hours of beam available at nominal intensity: 5.5
Hours of E832 data (special runs not included): 5.5
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: 1 kfrend.

========================================================================
Runs ended this shift:
Run 14092 - BEAM_832_KE3HOG (in progress)

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:
E832 Special low intensity run for ke3s. (SEM=2E12)

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

14092 5:34 Linac problems - no estimate.

6:02 Estimate 90 minutes for beam.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Occasional ETOT temp alarms from SDAQ -- check with ETOT experts during
day shift.

Low intensity running will end when the current tapes fill up.
Switch NM4AN to positive polarity before starting next run with standard
BEAM_832 configuration. Initial intensity request will be 4.2E12.